# Knowledge-Centric Templatic Views of Documents

### v1

@@ Abstract
Authors seeking to communicate with broader audiences often compose their ideas about the same underlying knowledge in different documents and formats – for example, as slide decks, newsletters, reports, brochures, etc. Prior work in document generation has generally considered the creation of each separate format to be different a task, developing independent methods for generation and evaluation. This approach is suboptimal for the advancement of AI-supported content authoring from both research and application perspectives because it leads to fragmented learning processes, redundancy in models and methods, and disjointed evaluation. Thus, in our work, we consider each of these documents to be templatic views of the same underlying knowledge, and we aim to unify the generation and evaluation of these templatic views of documents. We begin by introducing an LLM-powered method to extract the most important information from an input document and represent this information in a structured format. We show that this unified representation can be used to generate multiple templatic views with no supervision and with very little guidance, improving over strong baselines. We additionally introduce a unified evaluation method that is template agnostic, and can be adapted to building document generators for heterogeneous downstream applications. Finally, we conduct a human evaluation, which shows that humans prefer 82% of the downstream documents generated with our method. Furthermore, the newly proposed evaluation metric correlates more highly with human judgement than prior metrics, while providing a unified evaluation method.
Figure 1: Visualization of our method. Given an input document, we prompt the LLM to generate a Structure Unified Representation (SURe). We can use the SURe to prompt the model to generate a templatic view of the input document. We then evaluate the generations using our unified evaluation.
@@ 1 Introduction
Sharing information in multiple formats is an important part of communication in a variety of domains, as it allows for knowledge to be disseminated to a wider audience. A product manager may need to create a requirements document, a product pitch deck, and a product announcement newsletter. Likewise, a person on the job market may create a resume, a cover letter, and a personal website. We consider these documents to be templatic views of the same underlying knowledge.
@@
This is also true for the scientific domain, in which researchers need to create documents in multiple formats to effectively disseminate their work, – such as through conference talks, social media posts, poster presentations, and non-technical blog posts. Sharing knowledge in multiple formats broadens its audience and can help bridge the information gap between domain experts, researchers in adjacent fields, and even the general public – leading to greater understanding, new collaboration and accelerated progress in the field. With the significant increase in the pace of publication in recent years Bornmann and Mutz (2014), it has become essential for scientists to engage with the community and share their work in more compact, efficient, or easier to digest formats.
Past work on document generation has focused on developing methods of generation and evaluation for each specific document type individually Fu et al. (2021); Qiang et al. (2016); Chandrasekaran et al. (2020). However, tailoring methods to each template is time consuming and limiting, resulting in methods that are usable only for each intended template that cannot be compared across template types with specialized evaluation metrics. Our goal is to unify the tasks of generating and evaluating templatic views of documents. By doing so, we aim to allow system designers to adapt to new document types, templates and domains easily and efficiently, and to allow researchers to compare approaches more flexibly and fairly.
We begin by setting a few desiderata for our system for templatic view generation; it must be (1) unified, (2) adaptable, and (3) require minimal supervision. In order to design such a system to generate each of these templates, the method must be able to create structure and follow style norms associated with each template. We thus introduce an LLM-powered approach to generate a Structure Unified Representation (SURe) that represents the most important knowledge of an input document in a structured format. We demonstrate that this SURe can be used with no supervision and with very little prompted guidance about the output formats, to generate multiple templatic views of a document.
Meanwhile, automatic evaluation is crucial for researchers and developers to iterate more quickly than they otherwise would be capable of by relying solely on human evaluation. However, because in the past each templatic view has been tackled independently, previous work has used different evaluation metrics for each task, such as BERTScore for summarization Zhang* et al. (2020), Slide Level ROUGE for slide decks Qiang et al. (2016), or even a custom-trained, model-based metric for posters Wang et al. (2015). In this paper, we introduce Template Agnostic Evaluation (TAE) – a unified precision-recall style method of automatic evaluation that is template agnostic. Importantly, this evaluation framework is highly customizable, allowing users to easily integrate existing text-based metrics from the literature into its formulation and tailor its function to their specific use case.
@@
We evaluate our unified approach to templatic view generation and evaluation on 3 types of documents: slides, posters, and blog posts. In each of these, the inputs are scientific articles and the outputs are expected to be documents of the corresponding types. We use the DOC2PPT Fu et al. (2021), Paper-Poster Qiang et al. (2016), and LongSumm Chandrasekaran et al. (2020) datasets, respectively. We show that our method of generating a Structured Unified Representation leads to improvements in performance across each tasks, demonstrating that LLMs are capable of generating templatic views with structure-augmented prompts, despite having no supervision. We additionally conduct a human evaluation to validate both our document generation method and evaluation metric. We show that humans prefer the output yielded by our method 82% of the time and that our evaluation metric correlates more highly with human preference than other popular metrics, such as ROUGE and BERTScore Lin (2004); Zhang* et al. (2020).
While the focus of this paper is the transformation of scientific documents, this choice was primarily driven by the availability of publicly available data. We note that nothing about our unified framework is domain specific, and we believe our approach for the creation and evaluation of templatic views of documents is easily adaptable to novel domains, and document types. More generally, we believe our work is a step towards the broader and more challenging goal of AI-assisted authoring support, and future research in reasoning over and generation of long documents.
@@ 2 Related Work
Large Language Models (LLMs), which are central to our approach, have shown impressive capabilities in a variety of tasks Radford et al. (2019); Brown et al. (2020). Based on the transformer architecture Vaswani et al. (2017), LLMs have shown emergent abilities in tasks such as arithmetic and question answering Wei et al. (2022a). Similar to chain of thought prompting Wei et al. (2022b) and content planning prompting Wang et al. (2023b), we show that by generating an intermediate Structured Unified Representation (SURe) of an input document can improve performance over simply prompting the model to generate the final document Additionally, past work has explored using knowledge graphs to provide structure to LLMs in order to improve performance Pan et al. (2024); Rosset et al. (2020); Lewis et al. (2020). Similarly, we show that the structure provided by the SURe improves the quality of the final document, while being more easily adapted to new domains than a knowledge graph.
As past work has considered generation of templatic views to be separate tasks, methods for automatic evaluation of different templatic views has also been considered independently. LongSumm, the shared task introduced by Chandrasekaran et al. (2020), uses ROUGE scores to evaluate model performance Lin (2004). Fu et al. (2021) introduced Slide Level ROUGE to evaluate slide generation, which is a variant of ROUGE that contains a penalty for the number of slides. Qiang et al. (2016) used a trained regressor to predict the size of panels in a poster. In the summarization space, many automatic metrics for evaluation have been introduced such as BERTScore Zhang* et al. (2020), UniEval Zhong et al. (2022), BARTScore Yuan et al. (2021), BLANC Vasilyev et al. (2020), and MoverScore Zhao et al. (2019). However, these metrics are intended for a simple input document-summary setup, and do not take into account factors that affect the quality of other types of documents, such as structure and ordering. Our work is the first to introduce template agnostic evaluation, allowing uniform comparison of performance across template types.
Prompt Function Prompt
Generate the SURe
"Given the input text, extract the document title and authors.
For each section in the given input text, extract the most important sentences.
[…]
Output:"
[…]
"Summarize the following input in a <TEMPLATE TYPE>style.
Style parameters: <STYLE PARAMETERS>
[…]
Output:"
Table 1: Prompts used to generate the SURe and final LaTeX document. The template is pictured in Figure 2.
@@ 3 Data
Before introducing our approach to unified generation and evaluation of templatic document views, we begin by describing the data we use in this paper.
There is no existing dataset that includes multiple views of a single document. Instead, we evaluate our unified method, described in §4, on 3 datasets: DOC2PPT, LongSumm, and Paper-Poster Fu et al. (2021); Chandrasekaran et al. (2020); Qiang et al. (2016). These datasets are chosen because they each involve generating a different view of a document. We note that our method is not specific to the scientific domain. However, the scientific domain is one of the few domains with abundantly available public data of multiple templatic views of documents. We leave exploration of other domains to future work. The three datasets and their associated generation tasks are described below.
@@
We use use the DOC2PPT dataset introduced by Fu et al. (2021). DOC2PPT is a dataset of 5.8K scientific papers in Computer Science and their respective slide decks. We randomly sample 1K examples from this dataset for evaluation. The slides are provided as an image for each slide. We use the Azure OCR tool to extract the text from each slide [^1] https://learn.microsoft.com/en-us/azure/ai-services/computer-vision/overview-ocr.
@@
We use the LongSumm dataset introduced by Chandrasekaran et al. (2020), which includes blog posts of scientific papers in the Computer Science domain. LongSumm was originally introduced as a shared task for the purposes of long form summarization of scientific documents. Although LongSumm includes a blind set of 22 papers, our goal is to develop both a unified method of generation and evaluation across template types. As we are unable to compute our custom evaluation metric, described in §5, we exclude the blind test data from our evaluation. Instead, we use their released train split as our evaluation data, since our method does not require any training or supervision. Of the 531 released blog posts, we were able to access 505 of their papers. The other 26 papers included broken links or were behind a paywall. Longsumm also includes a set of extractive summaries from Lev et al. (2019). We exclude these summaries from our evaluation data.
@@
We use the Paper-Poster dataset introduced by Qiang et al. (2016). Paper-Poster is a dataset of 85 papers in Computer Science and Biology, and their respective scientific posters. We exclude 2 examples containing corrupted PDFs from our evaluation. Although Qiang et al. (2016) uses the source files to extract the text of posters for evaluation, they only release the PDF formats. To preprocess the reference posters, we found that automatic tools to extract text from documents did not handle the visual layout of posters well, so we manually extracted the text of the posters in this dataset. Note that this process was only done to obtain metrics for evaluation, and that our unsupervised generation method is capable of creating target documents without the need for reference data, or this dataset-specific manual extraction step.
For all 3 datasets, we use the Azure Document Layout tool to extract the text of the input papers. [^2] https://learn.microsoft.com/en-us/azure/ai-services/document-intelligence/concept-layout?view=doc-intel-4.0.0
@@ 4 Unified LLM-powered Generation of Templatic Views
Our goal is to develop a method of generation that is: (1) unified, (2) adaptable, and (3) requires minimal supervision. We introduce a 2 step method: generate a Structured Unified Representation (SURe) and then generate the templatic view. By first generating a SURe, we are providing the model with a structured representation of the knowledge in the input document. This representation unifies the resulting templatic views, and can be used to generate any new template styles, achieving criteria (1) and (2). Additionally, this method is a zero shot method, meaning it requires no previously annotated data or training, achieving criteria (3). This process is visualized in Figure 1.
4.1 SURe Generation
@@
The SURe is a JSON representation of the original document, consisting of a structured layout of the most important parts of the input. We provide the document to the model along with a template of the SURe. The prompt and template used can be found in Table 1 and Figure 2, respectively. By first generating a SURe, we are providing structure to the input document, while representing only the most important information, making it easier for the LLM to parse than the full input document. While we experiment with representing the SURe as a JSON in this paper, we note that a JSON format isn’t inherent to the SURe. The SURe could be represented in any format that provides structure to the input text (e.g. XML or Markdown) and we leave exploration of different formats to future work.
4.2 Templatic View Generation
We then feed the generated SURe as input back into the LLM, asking the model to generate the final output document. We ask the LLM to represent the final output document as a LaTeX file, as LaTeX is capable of formatting multiple styles of documents, including slides, posters, and blog posts. The prompts for each step can be found in Table 1. For each templatic view, the prompt to generate the final LaTeX document takes a short description of the desired output, which we refer to as style parameters. For the tasks we evaluate in this paper, we use the following style parameters:
  • Blogs: “Blogs should include pargraphs introducing the topic, a summary of the input document, and important takeaways. The blog should be more readable to a general audience than the input document.”
By incorporating the use of style parameters, our method is adaptable to new templatic views; the user only needs to write a short description of the template style. This means that although we evaluate our method on the scientific domain, our method is not limited and could be generalized to other domains and template styles, such as enterprise-specific document types. We leave the exploration of other domains to future work.
Figure 2: Template provided to generate the SURe.
5 Template Agnostic Evaluation
As past work has treated the generation of different templatic views as separate tasks, the evaluation of these types has also been developed independently. Unfortunately, none of the previously used metrics scale well to other document types or more generally to a unified evaluation of templatic views. ROUGE Lin (2004); Chandrasekaran et al. (2020) is only designed to evaluate text against a text reference, and does not take structure and ordering of information into account; the same is true for other fuzzy text comparison metrics Zhang* et al. (2020); Zhong et al. (2022); Yuan et al. (2021). Slide-level ROUGE Fu et al. (2021) partly solves the problem with a length penalty but still does not account for distribution of information, nor is it easily extensible to formats other than slides. Finally using a trained regressor Qiang et al. (2016) as a metric is only possible with availability of human-labeled data, which is expensive and time-consuming to curate for new domains and applications.
@@ 5 Template Agnostic Evaluation
Thus, our goal is to develop a framework of automatic evaluation that is template agnostic. In order to generalize to multiple templates, we introduce the concept of panels. A panel is a unit of organization within a document type, for which the placement and ordering of the panel is important to the overall flow of information in the document.
@@
For example, we consider panels to be each slide in a slide deck and each section on a poster. We consider the entirety of a blog post to be a single panel. Although we test our method on the tasks of slide, blog, and poster generation, the concept of panels is not limited to these document types. For example, each post on a social media thread could be considered a panel, or each page on a website. By generalizing to the idea of panels, we can define an evaluation metric that is template agnostic.
@@
We aim to unify the evaluation of templatic views by adapting and integrating prior metrics into a template agnostic precision-recall framework, which we refer to as Template-Agnostic Evaluation (TAE). TAE is not a new individual metric, but rather a framework of evaluation allowing generalization to new templates. For example, although ROUGE is not designed to capture posters, TAE can be used with ROUGE to evaluate poster generation.
The general formulation of TAE is as follows:
@@
in which $Q_{P}$ is the precision measure of quality (§5.1), $O_{P}$ is the precision penalty for order (§5.2), and $L$ is the penalty for length (§5.3). Similarly $Q_{R}$ is the recall measure of quality and $O_{R}$ is the recall penalty for order. The length penalty $L$ is non-reflexive, as discussed in §5.3. We calculate separate precision and recall scores, allowing evaluators to decide which measure is most important to them, including the ability to calculate an overall F-measure score.

### v2

@@ Abstract
Authors seeking to communicate with broader audiences often share their ideas in various document formats, such as slide decks, newsletters, reports, and posters. Prior work on document generation has generally tackled the creation of each separate format to be a different task, leading to fragmented learning processes, redundancy in models and methods, and disjointed evaluation. We consider each of these documents as templatic views of the same underlying knowledge/content, and we aim to unify the generation and evaluation of these templatic views. We begin by showing that current LLMs are capable of generating various document formats with little to no supervision. Further, a simple augmentation involving a structured intermediate representation can improve performance, especially for smaller models. We then introduce a novel unified evaluation framework that can be adapted to measuring the quality of document generators for heterogeneous downstream applications. This evaluation is adaptable to a range of user defined criteria and application scenarios, obviating the need for task specific evaluation metrics. Finally, we conduct a human evaluation, which shows that people prefer 82% of the documents generated with our method, while correlating more highly with our unified evaluation framework than prior metrics in the literature.
Figure 1: Visualization of our method to unify the generation and evaluation of templatic views of documents. Given an input document, we prompt the LLM to generate an intermediate representation. We can use the representation to prompt the model to generate a templatic view of the input document. We then evaluate the generations using our unified evaluation framework. The LLM represented in the figure is the same model.
@@ 1 Introduction
Sharing information is vital for communication and discourse across domains, as it allows for knowledge to be disseminated to a wider audience. This is often done by users through documents in multiple formats that nevertheless share some underlying knowledge. A product manager may need to create a requirements spec, a product pitch deck, and an announcement newsletter for the same project. Likewise, a person on the job market may create a resume, a cover letter, and a personal website. We consider these documents to be templatic views of the same underlying knowledge.
@@
This is equally true for the scientific domain, in which researchers create documents in multiple formats to effectively communicate and showcase their work, – such as through academic papers, conference talks, social media posts, poster presentations, and non-technical blog posts. Sharing knowledge in multiple formats broadens the audience and can help bridge the information gap between domain experts, researchers in adjacent fields, and even the general public, leading to greater understanding, collaborations and accelerated progress Bornmann and Mutz (2014).
Past work on document generation has focused on developing generation and evaluation methods specific to a single document type Fu et al. (2021); Qiang et al. (2016); Chandrasekaran et al. (2020). Narrow, custom methods tailored to individual document types are, nevertheless, time consuming to engineer and manage over the long term. For example, in an enterprise setting, it’s common to have dozens of occupation- and task-specific documents, each with their own template.Additionally, specific trained methods require data that may be expensive to acquire, or even be unavailable entirely. Meanwhile, LLMs have recently shown great success in long document generation Radford et al. (2019); Brown et al. (2020), indicating that this fragmentation of methods may no longer be necessary. Thus, our goal is to unify methods for both generating and evaluating templatic views of documents, allowing system designers and engineers to manage and adapt to a range of document types and domains easily and efficiently.
We begin by showing that LLMs are capable of diverse, structured document generation, requiring very little instructional guidance to do so effectively. Additionally, a few minor augmentations to the prompt – such as a structured, intermediate representation, and simple stylistic descriptions – can further improve downstream performance, especially for smaller, less resource intensive models. These findings have important implications on the deployment and scaling of unified, real-world AI-assisted document authoring systems.
In similar vein, we then introduce Template Adaptable Evaluation (TAE), departing from prior work’s task specific evaluation methods Zhang* et al. (2020); Qiang et al. (2016); Wang et al. (2015). TAE is a unified precision-recall style framework for automatic evaluation that is highly customizable, allowing users to easily integrate existing text-based metrics from the literature into its formulation and tailor it to their specific use case.Additionally, this framework allows developers to compare performance across document types, without needing to develop an evaluation metric for each individual template.
@@
We evaluate our unified approach for templatic view generation and evaluation on 3 types of documents: slides, posters, and blog posts Fu et al. (2021); Qiang et al. (2016); Chandrasekaran et al. (2020). Our experiments demonstrate that using a structured intermediate representation leads to improvements in performance across tasks, with greater gains for smaller language language models. In our human evaluation to validate both our unified document generation method and evaluation metric, we show that annotators prefer the output yielded by the structure-aware generation process 82% of the time and that our evaluation metric correlates more highly with human preference than other popular metrics. We release our code [^1] https://github.com/microsoft/knowledge-centric-templatic-views to support future research.
@@ 2 Related Work
Large Language Models (LLMs), which are central to our approach, have shown impressive capabilities in a variety of tasks Radford et al. (2019); Brown et al. (2020). Based on the transformer architecture Vaswani et al. (2017), LLMs have shown emergent abilities in tasks such as arithmetic and question answering Wei et al. (2022a). Similar to chain of thought prompting Wei et al. (2022b) and content planning prompting Wang et al. (2023b), we show that by generating an intermediate representation of an input document can improve performance over simply prompting the model to generate the final document from the original input.
As past work has tackled generation of templatic views as separate tasks, methods for automatic evaluation of different document types is fragmented. LongSumm, the shared task introduced by Chandrasekaran et al. (2020), uses ROUGE to evaluate model performance Lin (2004). Fu et al. (2021) introduced Slide Level ROUGE to evaluate slide generation, a variant that contains a penalty for the number of slides. Qiang et al. (2016) used a trained regressor. For summarization, many automatic evaluation metrics have been introduced such as BERTScore Zhang* et al. (2020), UniEval Zhong et al. (2022), BARTScore Yuan et al. (2021), BLANC Vasilyev et al. (2020), and MoverScore Zhao et al. (2019). However, these metrics are intended for a simple input document-summary setup, and do not take into account factors that affect the quality of other types of documents (e.g. structure). Our work is the first to introduce template adaptable evaluation, allowing uniform comparison of performance across template types.
@@ 3 Data
We begin by describing the data used in this paper. There is no existing dataset that includes multiple views of a single document. Instead, we evaluate our unified method, described in §4, on 3 existing datasets: DOC2PPT, LongSumm, and Paper-Poster Fu et al. (2021); Chandrasekaran et al. (2020); Qiang et al. (2016). These datasets are chosen because they each involve generating a different view of a document. Although our method is not specific to the scientific domain, it is one of the few domains with abundantly available public data of multiple templatic views  [^2] We acknowledge that scientific writing does have structural regularities that may influence unified document generation. Due to the lack of other available datasets we leave exploration of other domains to future work.. The three datasets and their associated generation tasks are described below.
@@
We use use the DOC2PPT dataset Fu et al. (2021), which contains 5.8K scientific papers in Computer Science and their respective slide decks. As Fu et al. (2021) do not release data splits or code, we randomly sample 1K examples from this dataset for evaluation. The slides are provided as an image for each slide. We use the Azure OCR tool to extract the text from each slide [^3] https://learn.microsoft.com/en-us/azure/ai-services/computer-vision/overview-ocr.
@@
We use the LongSumm dataset Chandrasekaran et al. (2020), which includes blog posts of scientific papers in the Computer Science domain. Since our approach requires no training or supervision, we use the entire training split from Longsumm as our evaluation set. Of the 531 publicly released blog posts in this set, we could only access 505, with the other 26 including broken links or being behind a paywall.
Notably, while Longsumm includes a blind test set of 22 papers, this test set only consists of inputs without their reference outputs, thus making it impossible to compute our custom evaluation metric (see §5). In the interest of completeness and comparison to prior work, we do, however submit runs from our systems to the leader board and report the results of this blind test set in Appendix D.
@@
We use the Paper-Poster dataset Qiang et al. (2016), which consists of a dataset of 85 papers in Computer Science and Biology, and their respective scientific posters; two examples containing corrupted PDFs are excluded. Although Qiang et al. (2016) release data splits, they do not release code or results for comparison. Given the small size of the dataset, we use it in its entirety for more robust results. While the authors uses the source files to extract the text of posters for evaluation, they only release the PDF formats. To preprocess the reference posters, we found that automatic tools to extract text from documents did not handle the visual layout of posters well, so we manually extracted the text of the posters in this dataset. Note that this process was only done to obtain evaluation scores, and that our unsupervised generation method is capable of creating target documents without the need for reference data.
For all 3 datasets, we use the Azure Document Layout tool to extract the text of the input papers. [^4] https://learn.microsoft.com/en-us/azure/ai-services/document-intelligence/concept-layout
@@ 4 Unified LLM-powered Generation of Templatic Views
The most straightforward way to transform documents between templatic views using LLMs, is to simply prompt the system to generate the target view given the input. However, similar to chain of thought prompting Wei et al. (2022b), we hypothesize that first generating a structured, intermediate representation of an input document and then reasoning over that representation will result in better generations than directly prompting the model. Our goal is to evaluate the capabilites of LLMs to generate long, structured documents, and experiment with how structured prompting can improve performance. We experiment with a simple general two-step process: first generate an intermediate representation, then generate the templatic view. These steps are described in greater detail below, and the process is visualized in Figure 1.
Intermediate Representation Generation.
@@
In this work, we set the intermediate representation to be a JSON consisting of a structured layout of the most important parts of the input. We provide the input document to the model along with a template of the representation and prompt it to extract the most important information from the input document, and format it in the given JSON structure. The exact prompts and JSON structure can be found in Appendix §A. While our experiments use a JSON intermediate representation, note that other formats that provide structure to the input text could be employed (e.g. XML or Markdown). Rather than trying to optimize for the best representation format, our goal is to show that this chain of extraction approach along with structured augmentation to prompts can aid the quality of generations from LLMs. We leave exploration of different formats and other prompt optimization to future work.
Templatic View Generation.
We then feed the generated representation as input back into the LLM, prompting the model to generate the final output document, represented as a LaTeX document. For each templatic view, the prompt to generate the final LaTeX document takes a short description of the desired output, which we refer to as a style parameter. For example, the style parameter for slide generation is as follows: “Slides should include a title page. Following slides should contain an informative slide title and short, concise bullet points. Longer slides should be broken up into multiple slides.” The use of style parameters makes our method adaptable to new templatic views; the user only needs to write a short description of the template style. Both the generation of the intermediate representations and the final documents require little to no prompt engineering. The prompts and style parameters can be found in Appendix §A.
5 Template Adaptable Evaluation
@@
Prior work on document generation has treated the evaluation of different templatic views as separate tasks. Thus, our goal is to develop a framework of automatic evaluation that is template adaptable. This not only allows us to compare performance across diverse datasets, it also removes the requirement of designing and maintaining individual metrics for each template. In order to generalize to multiple templates, we introduce the concept of panels. A panel is a unit of organization within a document type, for which the placement and ordering of the panel is important to the overall flow of information in the document.
@@
For example, we consider panels to be each slide in a slide deck and each section on a poster. We consider the entirety of a blog post to be a single panel. Although we test our method on the tasks of slide, blog, and poster generation, the concept of panels is not limited to these document types. For example, each post on a social media thread could be considered a panel, or each page on a website.
@@
We aim to unify the evaluation of templatic views by integrating prior metrics into a template adaptable precision-recall framework, which we refer to as Template-Adaptable Evaluation (TAE). TAE is not a new individual metric, but rather an evaluation framework that allows generalization to new templates. For example, TAE can even be used with ROUGE to evaluate poster generation.
The general TAE formulation is as follows:
@@
in which $Q_{P}$ is the precision measure of quality (§5.1), $O_{P}$ is the precision penalty for order (§5.2), and $L$ is the non-reflexive penalty for length (§5.3). Similarly, $Q_{R}$ is the recall quality measure and $O_{R}$ is the recall penalty for order. The precision-recall formulation allows evaluators to decide which measure is most important to them, or calculate an overall F-measure score.